OpenStudy (anonymous):

A kicker kicks a football toward the opponent's goal line during a game. The ball begins its flight with an initial velocity of 90 feet per second when it is kicked at a height of 2 feet by the kicker. To the nearest foot, what maximum height will the ball reach? Use the function h(t)=-16t^2+70t+2

OpenStudy (maddy1251):

I think I can help ya out. :D

OpenStudy (maddy1251):

I made an oops, I think... h '(t) = v(t) = -32t + 90 When ball reaches maximum height, velocity is zero -32t + 90 = 0 32 t = 90 t = 90 /32 = 45 /16 substitute t = 45/16 in h(t) h(45/16) = -16(45/16)^2 + 90(45/16) + 2 = 128.56 ft I think this is the correct answer now?

OpenStudy (anonymous):

thanks but i already answered it i got 159 >.<